# Tricritical Kibble–Zurek Scaling in Rydberg Atom Ladders

### v1

@@ Abstract
The Kibble–Zurek (KZ) mechanism is being actively explored on quantum simulation platforms. In this work, we study the KZ scaling around tricritical phase transition points, with Rydberg atom ladders as a concrete incarnation. The criticality is of Ising and Potts type for two- and three-leg ladders, respectively. When slowly ramping across or near the tricritical point from the disordered phase to the ordered phase, we obtain universal power-law scaling in agreement with conventional KZ predictions. We study a “tangential” KZ ramping that both begins and ends in the disordered phase, a novel protocol enabled by the two-dimensional phase diagram. The tangential KZ directly reveals subleading critical exponents of the critical point. Finally, we explore the regime of intermediate-speed ramping and find a dynamical analog of the celebrated Zamolodchikov’s $c$-theorem. Practically, our work provides an immediately relevant protocol for current experiments to pinpoint the elusive tricritical points. More broadly, tangential and intermediate-speed rampings go beyond the conventional KZ paradigm and introduce new insights into critical quantum dynamics.
@@
The Kibble–Zurek (KZ) mechanism concerns dynamical scaling behavior near a critical point [1, 2, 3, 4]. Originally proposed in the context of classical phase transitions, its quantum counterpart has been extensively explored both within [5, 6, 7, 8, 9, 10, 11, 10, 12, 13] and beyond [14, 15, 16, 17, 18, 19] the Landau–Ginzburg–Wilson paradigm. With recent advances in quantum simulators, numerous experimental studies on artificial quantum platforms, such as superconducting qubits [20, 21], ion traps [22, 23, 24], and Rydberg atom arrays [25, 26, 27, 28], have emerged. The KZ ramping technique is particularly well-suited for probing critical phenomena for two complementary reasons. First, quantum simulators allow for versatile dynamical control and site-resolving measurements. Second, dynamic probing bypasses the challenge of adiabatic preparation of a gapless critical state [29], where only a finite coherence time is available in these platforms.
@@
Although the KZ mechanism has been studied across a wide range of phase transitions, relatively few investigations have focused on a specific type of criticality: tricriticality [30, 31, 32, 33, 34]. This phenomenon can occur at the intersection of first- and second-order phase transitions [35, 36, 37, 38]. The tricritical point holds theoretical and practical importance. For instance, space-time supersymmetry, originally proposed in high-energy physics, can emerge in the low-energy regime of various models [39, 40, 41, 42, 43, 44], with one of the most experimentally promising cases being the $1+1$-dimensional tricritical Ising model [45, 46, 47]. Tricritical points also play an important role in quantum annealing protocols, where ramping near these points can accelerate optimization processes [48, 49, 50]. Furthermore, the proximity to first-order phase transitions raises questions about the validity of standard KZ scaling [51], suggesting that new phenomena may emerge in this context.
@@
Figure 1: (a, b) Two- and three-leg ladders that exhibit tricritical Ising and Potts points, respectively. Atoms of different colors correspond to different atomic species. The geometry is fixed (i.e., $r_{1,2}$), while the Rabi frequency $\Omega$ and detuning $\Delta$ can be tuned. (c) The phase diagram and ramping protocols. The solid (dashed) blue line denotes the second (first) order Ising/Potts transition, and a tricritical point lies in between. In Protocol I, we keep $\Omega$ fixed and ramp $\Delta$ from the disordered phase to the symmetry breaking phase, both across and near the tricritical point. In Protocol II, we consider a tangential ramping direction. In Protocol III, we select a sweet-spot direction to explore intermediate-speed KZ ramping.
@@
In this Letter, we propose an experimentally accessible model based on Rydberg atom ladders [52, 53, 54, 55], which can be easily implemented in state-of-the-art experiments (see Fig. 1(a,b)). Figure 1(c) presents the ground-state phase diagram, controlled by the Rabi frequency and detuning, with the key feature being the tricritical point connecting 1st- and 2nd-order (Ising/Potts) transition lines. The tricritical Ising (TCI) and tricritical Potts (TCP) transitions [56, 57] are realized respectively in the two- and three-leg ladders. In the renormalization group (RG) sense, these tricritical fixed points have two symmetric relevant operators on them [58, 59], as we detail below. This introduces new features compared to the extensively studied critical points with only one symmetric relevant operator [5, 6, 7, 8, 10, 11, 12].
@@
Furthermore, since the tricriticality only occurs at a single point in the phase diagram, generically (i.e. with non-vanishing deviation) only intermediate-distance physics is controlled by the tricritical point, with long-distance behaviors astray. Indeed, this is a concrete ramification of the celebrated Zamolodchikov’s $c$-theorem [60]. This renders measuring critical exponents in conventional materials challenging, but the very task fits well with the capability of site-resolved imaging in quantum simulators, using dynamical finite-size scaling [9, 61, 10, 13, 21, 28]. We capitalize on this through the ramping Protocols I, II, and III, as shown in Fig. 1(c). Protocol I is of immediate experimental relevance. The conceptual novelty of Protocol II lies in the fact that, throughout the ramping, the system parameters are in the disordered phase. Protocol III, with a sweet-spot direction enabling a much wider range of ramping speeds, reveals an intriguing crossover behavior from tricriticality to its parent second-order criticality, where a dynamical analog of Zamolodchikov’s $c$-theorem is proposed.
@@
The model.— The microscopic models, two- and three-leg Rydberg atom ladders, are modified from a previous model [47]. In both the original and the current models, the system consists of two species of atoms [62, 63, 64, 65], assembled in an alternative pattern along the ladder, see Fig. 1(a, b). Each atom is modeled by a two-state system, with the ground state $|g\rangle$ and the Rydberg state $|r\rangle$. A laser couples both states and determines the Rabi frequency $\Omega$ and detuning $\Delta$ through the laser strength and frequency, respectively. The two species are chosen such that Rydberg atoms of the same species repel, while those of different species attract. The Hamiltonian reads
@@
where $X,Y$ denote the species of the atoms $i$ and $j$, and $r_{ij}$ is the inter-atom distance. $n=|r\rangle\langle r|$ is a projector onto the Rydberg state or equivalently the Rydberg number operator. $\sigma_{x}=|g\rangle\langle r|+\mathrm{h.c.}$ couples the ground state and the Rydberg state. $r_{2}$ is chosen to be reasonably small, so that only one atom within a rung can be excited to the Rydberg state due to Rydberg blockade [66].
@@
In the original model, one fixes $r_{2}$ and $\Omega$, and tunes two parameters, namely $r_{1}$ and $\Delta$. In the current work, we consider instead tuning $\Omega$ and $\Delta$, keeping the geometry fixed throughout, see Fig. 1(c). In fact, the phase diagram of the original model has the same structure as in Fig. 1(c) (with $\Omega$ replaced by $r_{1}$) thanks to universality. However, the new setup has the virtue that both $\Omega$ and $\Delta$ are determined by the laser, and are experimentally readily tuned in a dynamical fashion. In contrast, the lattice geometry is unchangeable during the dynamical evolution  [^1] Indeed, the atoms are arranged to a desired geometry by using optical tweezers. During the evolution, with atoms excited to Rydberg states, the tweezers could easily ionize the Rydberg atoms, causing significant atom loss.. The upshot is that, in the original model, within each experimental run, one can only move in the horizontal (i.e. detuning) direction in the phase diagram. In the current setup, on the other hand, one is able to move along any path in the phase diagram.
@@
The ladder models have a $\mathbb{Z}_{2}$ ($S_{3}$) symmetry that permutes the two (three) legs. For small $\Delta$, the atoms remain in the ground state, and the system is in the disordered phase. For large $\Delta$, the atoms prefer to be excited to the Rydberg state. However, with Rydberg blockade in action, the best one can do is to excite one of the two (three) legs. This breaks the symmetry to $\varnothing$ (reflective $\mathbb{Z}_{2}$), and hence we expect the transition to be Ising (three-state Potts) type. For small $\Omega$, the Hamiltonian is dominated by diagonal terms in the $Z$ basis, and the transition tends to be first-order. For large $\Omega$, the quantum fluctuation dominates, and the transition tends to be second-order. In between, TCI (TCP) points emerge, with distinct critical exponents compared to the Ising (Potts) case. All these predictions are corroborated by numerical simulations using density matrix renormalization group (DMRG). With the model and the phase diagram at hand, we next turn to dynamical evolutions. In the main text, we focus on the two-leg ladder. For the three-leg case, results repeat almost verbatim and we postpone them to the End Matter.
Figure 2: (a) Binder ratio $U$ as function of $g=\Delta-\Delta_{\mathrm{TCI}}\propto t$ for various ramping speed $s$. The first peak is marked with a cross. The ground state Binder ratio (black dashed line) is included for reference. (b) KZ ramping across the TCI point, with data collapse achieved for $s^{\mu}L$. (c) KZ ramping near the TCI point, with data collapse achieved for $hL^{1/\nu^{\prime}}$. (d) For tangential KZ, we achieve data collapse with $s^{\mu^{\prime}}L$. In the insets, the data collapse figure of merit $\mathcal{D}$ is plotted against the corresponding critical exponents.
@@
KZ across and near the TCI point.— To begin with, we consider ramping from the disordered phase to the ordered phase, similar to the conventional KZ protocol. This is shown in Fig. 1(c) as Protocol I, where the ramping parameter is given by $g=\Delta-\Delta_{\mathrm{TCI}}$. The ground-state correlation length diverges as $\xi_{\mathrm{GS}}\sim g^{-\nu}$, where $\nu$ is the critical exponent. Since the correlation length $\xi$ itself has length dimension, i.e., $[\xi]=1$, the dimension of the ramping parameter is $[g]=-1/\nu$. Next, we define the dynamical process: ramping $g$ with speed $s$, i.e., $g=st$, with $t=0$ set as the time passing the phase boundary. By this, one can recover the celebrated KZ scaling of the correlation length, $\xi_{\mathrm{KZ}}\sim s^{-\mu}$, by noting the dimension of speed: $[s]=[g]-[t]=-1/\nu-z\equiv-1/\mu$, where $z$ is the dynamical critical exponent.
@@
In the quantum simulators we consider, however, the systems are finite-sized, meaning that the correlation length is limited at the critical point rather than diverging. This leads to a deviation from the $\xi_{\mathrm{KZ}}\propto s^{-\mu}$ scaling. Fortunately, as in the equilibrium case, finite-size scaling also applies to the KZ mechanism, namely $\xi_{\mathrm{KZ}}/L=\mathcal{F}(s^{\mu}L)$ [9, 61, 10, 13, 21, 28], where $L$ is the system size. However, extracting the length scale $\xi_{\mathrm{KZ}}$ from the two-point correlation function introduces ambiguity, and this long-distance feature can easily be affected by noise [28]. Therefore, we choose the Binder ratio $U$ to characterize the dynamics [68]. The main advantages of $U$ are twofold. First, it is dimensionless and does not scale by itself, thus eliminating one additional scaling parameter to be determined. Second, unlike the correlation length, it is a number defined without the need of fitting. Specifically, the time-dependent Binder ratio is defined through the order parameter $M$ as
@@
The order parameter is defined as $M=\sum_{i}n_{ia}-n_{ib}$, where $a$ and $b$ denote the atoms within each rung, and $|\psi(t)\rangle$ is the wavefunction evolved by the Hamiltonian (1) up to time $t$.
@@
Figure 2(a) shows how the Binder ratio changes over time when ramping across the critical point at different speeds $s$. We observe that the Binder ratio oscillates rather than directly reaching a plateau, so we need to be cautious when using Binder data to determine scaling laws. The Binder curves at different speeds should show similar behavior due to the scaling invariance of the criticality. We choose the first peak after crossing the critical point ($t>0$) as a reference point for finite-size scaling. This is known as finite-time scaling [69, 70, 71], meaning that we do not need to wait for a long time for the quantity to saturate; instead, we use data that is relatively close to the critical point. This is crucial for tricritical physics, as data far from the critical point may not obey tricriticality, as we will see later.
@@
We now discuss our first ramping scheme, Protocol I, which involves ramping across or near the tricritical point, as shown in Fig. 1(c). Experimentally, this corresponds to ramping the detuning $\Delta$ within a range of fixed Rabi frequency $\Omega$ around $\Omega_{\mathrm{TCI}}$, with the deviation from the tricritical point parametrized by $h=\Omega-\Omega_{\mathrm{TCI}}$. While the tricritical point $\Omega_{\mathrm{TCI}}$ can be determined with high precision numerically [66], achieving this can be tricky experimentally. We will see that without knowing $\Omega_{\mathrm{TCI}}$ a priori, using scaling techniques we can still extract the critical exponents. A side product of our protocol is a method of finding $\Omega_{\mathrm{TCI}}$.
@@
From a theoretical perspective, scanning near the tricriticality provides information about two relevant perturbations, resulting in both leading and subleading critical exponents [72, 73]. In particular, the leading critical exponent, $\mu$, can be extracted by ramping the parameter $g$ at speed $s$, following the standard KZ scaling. Simultaneously, the parameter $h$ is associated with a subleading critical exponent, $\nu^{\prime}$, with $[h]=-1/\nu^{\prime}$ (cf. $[g]=-1/\nu$). Note that $h^{\nu^{\prime}}L$ forms a dimensionless combination. These critical exponents can be obtained from conformal field theory (CFT), summarized in Table 1.
@@
universality class $c$ $\nu$ $\mu$ $\nu^{\prime}$ $\mu^{\prime}$
two-leg       Ising 1/2 1 1/2 / /
      TCI 7/10 5/9 5/14 5/4 5/9
three-leg    Potts 4/5 5/6 5/11 / /
       TCP 6/7 7/12 7/19 7/4 7/11
Table 1: The central charge and critical exponents of the universality classes [58]. The dynamical exponent $z$ for all cases is 1, due to the emergent Lorentzian symmetry.
Figure 3: (a) RG flow near the TCI. (b, c) Ramping with a broad range of ramping speed $s$. Both figures are plotted with the same data but with different scalings of $s$. (b) For slower ramping (upper right region), the TCI critical exponent allow the best data collapse. (c) For faster ramping (lower left region), the Ising critical exponent allow the best data collapse.
We perform extensive numerical simulations of the dynamical evolution using the time-evolving block decimation (TEBD) algorithm with open boundary conditions [74, 75, 25, 76, 28]. Using the Binder peaks extracted from Fig. 2(a) for different system sizes $L$, ramping speeds $s$, and ramping paths parameterized by $h$, we verify Eq. (3). In particular, we expect that for exponents $\mu$ and $\nu^{\prime}$ that match the universal ones, $U_{\mathrm{peak}}$ as a function of $s^{\mu}L$ and $h^{\nu^{\prime}}L$ will exhibit an optimal data collapse. Our numerical results lend full support for this expectation, as shown in the main figures of Fig. 2(b, c). Figure 2(b) corresponds to $h=0$, i.e., ramping across the tricritical point, and we fix $s^{\mu}L$ and collect data for a window of $\Omega$ in Fig. 2(c). In the inset of Fig. 2(b), we test different values of $\mu$ by a figure of merit $\mathcal{D}=\sqrt{\sum_{i}d_{i}^{2}}$, where $d_{i}$ is the vertical distance between the data point at $i$ and the smooth fitting curve value at the same horizontal value of $s^{\mu}L$. We plot $\mathcal{D}$ as a function of $\mu$, with the minimum corresponding to the optimal data collapse, which agrees closely with the CFT prediction. Next, in the inset of Fig. 2(c), we show that we can obtain $\nu^{\prime}$ even without the knowledge of $\Omega_{\mathrm{TCI}}$ by taking the latter as another optimization parameter. The contour of the figure of merit shows a good agreement with both $\nu^{\prime}$ and the earlier-used $\Omega_{\mathrm{TCI}}$.
@@
Tangential KZ scaling.— The two-dimensional phase diagram with a curved phase boundary allows a novel type of KZ scaling, which we dub “tangential” to highlight how the dynamical path touches the phase boundary. The conceptual novelty here lies in the fact that throughout the ramping one stays exclusively in the disordered phase and never enters the ordered phase, unlike previous KZ protocols. This fact leaves uncertain some heuristic arguments commonly found in the literature, where domain walls or generally topological defects that are only well-defined in the symmetry-breaking phase play a key role. Nevertheless, scaling relations that do not depend on detailed modeling of the dynamics remain valid.
In the tangential ramping, we are effectively ramping the subleading operator, with the leading operator vanishing. Hence the scaling relation is modified to

### v2

@@ Abstract
The Kibble–Zurek (KZ) mechanism has been extensively studied in various second-order phase transitions, yet the case of tricriticality—the point where second-order phase transition lines terminate—remains experimentally elusive. Here, we theoretically propose probing KZ scaling at tricritical points using Rydberg atom arrays arranged as two- and three-leg ladders, which realize the tricritical Ising and tricritical Potts universality classes. By slowly ramping the Rabi frequency and detuning, we extract two relevant tricritical exponents, $\nu$ and $\nu^{\prime}$, both via conventional paths from the disordered to the ordered phase and via “tangential” paths confined entirely within the disordered phase. At faster speeds, ramping dynamics go beyond the standard KZ paradigm: data collapse analysis using the parent critical exponents (rather than the tricritical ones) reveals renormalization group flows toward the adjacent second-order critical line, and we identify it as a dynamical analog of Zamolodchikov’s $c$-theorem. Our protocol is readily implementable on existing Rydberg quantum simulators. This provides a direct route to measuring distinct tricritical exponents which can reveal an emergent spacetime supersymmetry constraint $1/\nu-1/\nu^{\prime}=1$. Moreover, this work deepens our theoretical understanding and opens new avenues for exploring beyond-KZ quantum dynamics with rich renormalization group structure.
@@
The Kibble–Zurek (KZ) mechanism concerns dynamical scaling behavior near a critical point [kibble1976topology, zurek1985cosmological]. Originally proposed in the context of classical phase transitions, its quantum counterpart has been extensively explored both within [zurek2005dynamics, dziarmaga2005dynamics, polkovnikov2005universal, polkovnikov2011colloquium, Sandvik2011, kolodrubetz2012nonequilibrium, Sondhi2012, kolodrubetz2012nonequilibrium, del2014universality, rossini2021coherent] and beyond [Bermudez2009Topology, bermudez2010dynamical, ChenInsulatorKZ2020, Nie2024topo, shu2025equilibration, Deng2025] the Landau–Ginzburg–Wilson paradigm. With recent advances in quantum simulators, numerous experimental studies on artificial quantum platforms, such as superconducting qubits [king2022coherent, king2023quantum], ion traps [ulm2013observation, pyka2013topological, li2023probing], and Rydberg atom arrays [keesling2019quantum, ebadi2021quantum, manovitz2025quantum, zhang2025near], have emerged. The KZ ramping technique is particularly well-suited for probing critical phenomena for two complementary reasons. First, quantum simulators allow for versatile dynamical control and site-resolving measurements. Second, dynamical probing bypasses the challenge of adiabatic preparation of a gapless critical state [fang2024probing], where only a finite coherence time is available on these platforms.
@@
At the intersection of first- and second-order phase transitions, a particular phenomenon known as tricriticality occurs. The tricritical point holds theoretical and practical importance. For instance, spacetime supersymmetry, originally proposed in high-energy physics, can emerge in the low-energy regime of various models [grover2014emergent, jian2015susy, jian2017susy, li2018numerical, Li2020TCI, zeng2024nonequilibrium], with one of the most experimentally promising cases being the $1+1$-dimensional tricritical Ising model [Friedan1985, Zamolodchikov1986multicritical, li2024supersymmetry, cheng2025schwinger]. Tricritical points also play an important role in quantum annealing protocols, where ramping near these points can accelerate optimization processes [wang2022mbl, zhang2024cyclic, zhang2025computational]. Furthermore, the proximity to first-order phase transitions raises questions about the validity of standard KZ scaling [suzuki2024topological], suggesting that new phenomena may emerge in this context. Despite its importance, no experimentally feasible protocol of tricritical KZ effect has been proposed [Mukherjee2007XY, divakaran2009defect, deng2009anomalous, mukherjee2010adiabatic, patra2011non].
@@
Figure 1: Rydberg ladders and Kibble-Zurek protocols. a, b Two- and three-leg ladders that exhibit tricritical Ising and Potts points, respectively. Atoms of different colors correspond to different atomic species. The geometry is fixed (i.e., $r_{1,2}$), while the Rabi frequency $\Omega$ and detuning $\Delta$ can be tuned. c The schematic phase diagram and ramping protocols. The solid (dashed) blue line denotes the second (first) order Ising/Potts transition, and a tricritical point lies in between. In Protocol I, we keep $\Omega$ fixed and ramp $\Delta$ from the disordered phase to the symmetry breaking phase, both across and near the tricritical point. In Protocol II, we consider a tangential ramping direction. The orange dot arrow refers to ramping of higher speeds, where a dynamical analog to Zamolodchikov’s $c$-theorem is proposed.
@@
Here, we propose an experimentally accessible model based on Rydberg atom ladders [Sarkar2023ladder, zhang2025ladder, kerschbaumer2024ladder, Chepiga2025ladder], as illustrated in Fig. 1a and b. Figure 1c presents the ground-state phase diagram, controlled by the Rabi frequency and detuning, with the key feature being the tricritical point connecting 1st- and 2nd-order (Ising/Potts) transition lines. The tricritical Ising (TCI) and tricritical Potts (TCP) transitions [Nienhuis1979TCP] are realized respectively in the two- and three-leg ladders. From the perspective of renormalization group (RG) theory, these tricritical fixed points are characterized by two symmetry-preserving relevant operators [DiFrancescoCFTBook, CardyBook], introducing rich RG flow behaviors. These additional operators significantly impact the KZ dynamics, in contrast to the more widely studied critical points, which feature only a single symmetry-preserving relevant operator.
@@
Furthermore, since the tricriticality only occurs at a single point in the phase diagram, generically (i.e. with non-vanishing deviation) only intermediate-distance physics is controlled by the tricritical point, with long-distance physics exhibiting more conventional behavior. Indeed, this is a concrete consequence of the celebrated Zamolodchikov’s $c$-theorem [zamolodchikov1986irreversibility]. Measuring tricritical exponents is therefore challenging in conventional materials, but this very task fits well with the capability of site-resolved imaging in quantum simulators, using dynamical finite-size scaling [Sandvik2011, liu2014dynamic, kolodrubetz2012nonequilibrium, rossini2021coherent, king2023quantum, zhang2025near]. We capitalize on this through the ramping Protocols I and II, as shown in Fig. 1c. Protocol I is of immediate experimental relevance. The conceptual novelty of Protocol II lies in the fact that, throughout the ramping, the system parameters are in the disordered phase, which is beyond the framework relating KZ to domain wall formation. These protocols provide a direct route to measuring distinct tricritical exponents, which can reveal an emergent spacetime supersymmetry constraint in the two-leg case. Moreover, by exploring a much wider range of ramping speeds, the results reveal intriguing crossover behavior from tricriticality to its parent second-order criticality, where a dynamical analog of Zamolodchikov’s $c$-theorem is proposed.
Results
Rydberg models and equilibrium phase diagrams
@@ Results
The microscopic models, two- and three-leg Rydberg atom ladders, are modified from a previous model [li2024supersymmetry]. In both the original and the current models, the system consists of two species of atoms [Zhan2017PRL, Zhan2022PRL, Hannes2022PRX, anand2024dual], assembled in an alternating pattern along the ladder, see Fig. 1a, b. Each atom is modeled by a two-state system, with the ground state $|g\rangle$ and the Rydberg state $|r\rangle$. A laser couples both states and determines the Rabi frequency $\Omega$ and detuning $\Delta$ through the laser strength and frequency, respectively. The two species are chosen such that Rydberg atoms of the same species repel, while those of different species attract. The Hamiltonian reads
@@
where $X,Y$ denote the species of the atoms $i$ and $j$, and $r_{ij}$ is the inter-atom distance. $n=|r\rangle\langle r|$ is a projector onto the Rydberg state or, equivalently, the Rydberg number operator. $\sigma_{x}=|g\rangle\langle r|+\mathrm{h.c.}$ couples the ground state and the Rydberg state. $r_{2}$ is chosen to be reasonably small, so that only one atom within a rung can be excited to the Rydberg state due to Rydberg blockade.
@@
In the original model, one fixes $r_{2}$ and $\Omega$, and tunes two parameters, namely $r_{1}$ and $\Delta$. In the current work, we consider instead tuning $\Omega$ and $\Delta$, keeping the geometry fixed throughout, see Fig. 1c. In fact, the phase diagram of the original model has the same structure as in Fig. 1c (with $\Omega$ replaced by $r_{1}$) thanks to universality (see also the numerical phase diagram in Supplementary Note 1). However, the new setup has the virtue that both $\Omega$ and $\Delta$ are determined by the laser, and are experimentally readily tuned in a dynamical fashion. In contrast, the lattice geometry is unchangeable during the dynamical evolution. Indeed, the atoms are arranged in a desired geometry by using optical tweezers. During the evolution, with atoms excited to Rydberg states, the tweezers could easily ionize the Rydberg atoms, causing significant atom loss. The upshot is that, in the original model, within each experimental run, one can only move in the horizontal (i.e. detuning) direction in the phase diagram. In the current setup, on the other hand, one is able to move along any path in the phase diagram.
@@
The ladder models have a $\mathbb{Z}_{2}$ ($S_{3}$) symmetry that permutes the two (three) legs. For small $\Delta$, the atoms remain in the ground state, and the system is in the disordered phase. For large $\Delta$, the atoms prefer to be excited to the Rydberg state. However, with Rydberg blockade in action, the best one can do is to excite one of the two (three) legs. This breaks the symmetry to $\varnothing$ (reflective $\mathbb{Z}_{2}$), and hence we expect the transition to be Ising (three-state Potts) type. For small $\Omega$, the Hamiltonian is dominated by diagonal terms in the $Z$ basis, and the transition tends to be first-order. For large $\Omega$, the quantum fluctuation dominates, and the transition tends to be second-order. In between, TCI (TCP) points emerge, with distinct critical exponents compared to the Ising (Potts) case. All these predictions are corroborated by numerical simulations using density matrix renormalization group (DMRG) [White1992, MPS_2011], see Methods for details. With the model and the phase diagram at hand, we next turn to dynamical evolutions. In the main text, we focus on the two-leg ladder. For the three-leg case, results repeat almost verbatim and we postpone them to the Supplementary Note 2.
KZ scalings across and near TCI point
@@
To begin with, we consider ramping from the disordered phase to the ordered phase, similar to the conventional KZ protocol. This is shown in Fig. 1c as Protocol I, where the ramping parameter is given by $g=\Delta-\Delta_{\mathrm{TCI}}$. The ground-state correlation length diverges as $\xi_{\mathrm{GS}}\sim|g|^{-\nu}$, where $\nu$ is the critical exponent. Since the correlation length $\xi$ itself has length dimension, i.e., $\xi]=-1$, the dimension of the ramping parameter is $g]=1/\nu$. Next, we define the dynamical process: ramping $g$ with speed $s$, i.e., $g=st$, with $t=0$ set as the time passing the phase boundary. By this, one can recover the celebrated KZ scaling of the correlation length, $\xi_{\mathrm{KZ}}\sim s^{-\mu}$, by noting the dimension of speed: $s]=[g]-[t]=1/\nu+z\equiv 1/\mu$, where $z$ is the dynamical critical exponent.
@@
In the quantum simulators we consider, however, the systems are finite-sized, meaning that the correlation length is limited at the critical point rather than diverging. This leads to a deviation from the $\xi_{\mathrm{KZ}}\propto s^{-\mu}$ scaling. Fortunately, as in the equilibrium case, finite-size scaling also applies to the KZ mechanism, namely $\xi_{\mathrm{KZ}}/L=\mathcal{F}(s^{\mu}L)$ [Sandvik2011, liu2014dynamic, kolodrubetz2012nonequilibrium, rossini2021coherent, king2023quantum, zhang2025near], where $L$ is the linear system size (e.g., the length of the ladder in our quasi-1D geometry). However, extracting the length scale $\xi_{\mathrm{KZ}}$ from the two-point correlation function introduces ambiguity, and this long-distance feature can easily be affected by noise [zhang2025near]. Therefore, we choose the Binder ratio $U$ to characterize the dynamics. The main advantages of $U$ are twofold. First, it is dimensionless and does not scale by itself, thus eliminating one additional scaling parameter to be determined. Second, unlike the correlation length, it is a number defined without the need of fitting. Specifically, the time-dependent Binder ratio is defined through the order parameter $M$ as
@@
The order parameter is defined as $M=\sum_{i}n_{ia}-n_{ib}$, where $a$ and $b$ denote the atoms within each rung, and $|\psi(t)\rangle$ is the wavefunction evolved by the Hamiltonian (1) up to time $t$. This quantity can be readily measured in Rydberg platforms using site-resolved in situ imaging.
@@
Figure 2a shows how the Binder ratio changes over time when ramping across the critical point at different speeds $s$. We observe that the Binder ratio oscillates rather than directly reaching a plateau, so we need to be cautious when using Binder data to determine scaling laws. The Binder curves at different speeds should show similar behavior due to the scaling invariance of the criticality. We choose the first peak after crossing the critical point ($t>0$) as a reference point for finite-size scaling. This is known as finite-time scaling [gong2010finite, Huang2014FTS, shu2025emergent], meaning that we do not need to wait for a long time for the quantity to saturate; instead, we use data that is relatively close to the critical point. This is crucial for tricritical physics, as data far from the critical point may not obey tricriticality, as we will see later.
Figure 2: KZ scaling of Binder ratio in the TCI case. TEBD simulations using Hamiltonian (1) with $N=2L$ atoms and $1/r^{6}$ interactions for (ad). a Binder ratio $U$ as function of $g=\Delta-\Delta_{\mathrm{TCI}}\propto t$ for various ramping speed $s$ with $L=16$. The first peak is marked with a cross. The ground state Binder ratio for open boundary conditions (black dashed line) is included for reference. b KZ ramping across the TCI point, with data collapse achieved for $s^{\mu}L$. c KZ ramping near the TCI point, with data collapse achieved for $hL^{1/\nu^{\prime}}$. d For tangential KZ, we achieve data collapse with $s^{\mu^{\prime}}L$. In the insets, the data collapse figure of merit $\mathcal{D}$ is plotted against the corresponding critical exponents. eh show analogous results using an effective spin-1 model with size $L$ for TEBD simulations. The gray dashed curves represent quadratic fits in (b,c,f,g) and cubic fits in (d,h).
@@
We now discuss our first ramping scheme, Protocol I, which involves ramping across or near the tricritical point, as shown in Fig. 1c. Experimentally, this corresponds to ramping the detuning $\Delta$ within a range of fixed Rabi frequency $\Omega$ around $\Omega_{\mathrm{TCI}}$, with the deviation from the tricritical point parametrized by $h=\Omega-\Omega_{\mathrm{TCI}}$. While the tricritical point $\Omega_{\mathrm{TCI}}$ can be determined with high precision numerically, achieving this can be tricky experimentally. We will see that without knowing $\Omega_{\mathrm{TCI}}$ a priori, using scaling techniques we can still extract the critical exponents. A side product of our protocol is a method of finding $\Omega_{\mathrm{TCI}}$.
@@
From a theoretical perspective, scanning near the tricriticality provides information about two relevant perturbations, resulting in both leading and subleading critical exponents [Wilding1996, Kwak2015, Mathey2020, Ladewig2020]. In particular, the leading critical exponent, $\mu$, can be extracted by ramping the parameter $g$ at speed $s$, following the standard KZ scaling. Simultaneously, the parameter $h$ is associated with a subleading critical exponent, $\nu^{\prime}$, with $h]=1/\nu^{\prime}$ (see Methods and Fig. 4 for more details). Note that $h^{\nu^{\prime}}L$ forms a dimensionless combination. These critical exponents can be obtained from conformal field theory (CFT), summarized in Table 1.
Universality class $c$ $\nu$ $\mu$ $\nu^{\prime}$ $\mu^{\prime}$
two-leg Ising $1/2$ 1 $1/2$ / /
two-leg TCI $7/10$ $5/9$ $5/14$ $5/4$ $5/9$
three-leg Potts $4/5$ $5/6$ $5/11$ / /
three-leg TCP $6/7$ $7/12$ $7/19$ $7/4$ $7/11$
Table 1: The central charge and critical exponents of the universality classes. The dynamical exponent $z$ for all cases is 11, due to the emergent Lorentz symmetry.
@@
We perform extensive numerical simulations of the dynamical evolution using the time-evolving block decimation (TEBD) algorithm with open boundary conditions [TEBD, Verstraete2004TEBD, keesling2019quantum, Chepiga2024KZ, zhang2025near]. Using the Binder peaks extracted from Fig. 2a for different system sizes $L$, ramping speeds $s$, and ramping paths parameterized by $h$, we verify Eq. (3). In particular, we expect that for exponents $\mu$ and $\nu^{\prime}$ that match the universal ones, $U_{\mathrm{peak}}$ as a function of $s^{\mu}L$ and $h^{\nu^{\prime}}L$ will exhibit an optimal data collapse. Our numerical results lend full support for this expectation, as shown in the main figures of Fig. 2b, c. Figure 2b corresponds to $h=0$, i.e., ramping across the tricritical point, and we fix $s^{\mu}L$ and collect data for a window of $\Omega$ in Fig. 2c. In the inset of Fig. 2b, we test different values of $\mu$ by a figure of merit $\mathcal{D}=\sqrt{\sum_{i}d_{i}^{2}}$, where $d_{i}$ is the vertical distance between the data point at $i$ and the smooth fitting curve value at the same horizontal value of $s^{\mu}L$. We plot $\mathcal{D}$ as a function of $\mu$, with the minimum corresponding to the optimal data collapse, which agrees closely with the CFT prediction. Next, in the inset of Fig. 2c, we show that we can obtain $\nu^{\prime}$ even without the knowledge of $\Omega_{\mathrm{TCI}}$ by taking the latter as another optimization parameter. The contour of the figure of merit shows a good agreement with both $\nu^{\prime}$ and the earlier used $\Omega_{\mathrm{TCI}}$, and $\delta\Omega_{\mathrm{TCI}}$ in the inset of Fig. 2c represents the difference from the value obtained from DMRG.
Additionally, we perform larger-scale simulations by making certain approximations. Since $r_{2}$ is sufficiently small, two atoms in a rung cannot be simultaneously excited to the Rydberg state. Hence, each rung can be recognized as a spin-1 mode (see the Methods section for more details). With such a spin-1 representation, we can simulate much larger system sizes, more than double the original sizes, as shown in Fig. 2f–h. The conclusions drawn from the full $N=2L$ atom simulations remain unchanged, achieving the same universal scaling behavior.
@@
Figure 3: RG perspective of KZ in the TCI case. a RG flow (black lines) near the TCI. The blue line indicates the numerical phase boundary determined by Binder ratio and central charge (see Methods). b, c Ramping with a broad range of ramping speed $s$. Both figures are plotted with the same data but with different scalings of $s$. For slower ramping (upper right region in (b)), the TCI critical exponent allow the best data collapse. For faster ramping (lower left region in (c)), the Ising critical exponent allow the best data collapse.
Tangential KZ scaling
The two-dimensional phase diagram with a curved phase boundary allows a novel type of KZ scaling, which we dub “tangential” to highlight how the dynamical path touches the phase boundary. The conceptual novelty here lies in the fact that throughout the ramping one stays exclusively in the disordered phase and never enters the ordered phase, unlike previous KZ protocols. This fact leaves uncertain some heuristic arguments commonly found in the literature, where domain walls or generally topological defects that are only well defined in the symmetry-breaking phase play a key role. Nevertheless, scaling relations that do not depend on detailed modeling of the dynamics remain valid.
In the tangential ramping, we are effectively ramping the subleading operator, with the leading operator vanishing (see Methods for more details). Hence, the scaling relation is modified to